# H-GAR: A Hierarchical Interaction Framework via Goal-Driven Observation-Action Refinement for Robotic Manipulation

### v1

@@ 1  Introduction
Effective planning and manipulation in robotics (chatvla; cogact; dexgraspvla; autoeval; gr1; gr2; roboagent; zhu2025emosym; zhu2025uniemo; li2025lion; xie2025gui; shao2024detecting) require the ability to anticipate both how the environment evolves and how actions unfold over time. To this end, a growing body of research (cot; liang2024dreamitate; li2025unified) has shown that jointly predicting future visual observations and action sequences brings substantial benefits to robotic manipulation. This owes to the fact that anticipated visual observations provide rich contextual cues, including spatial layouts, object affordances, and interaction dynamics, which help guide action planning. In parallel, predicted actions expose the causal structure of interactions, enabling more accurate forecasting of how scenes evolve over time. This bidirectional coupling between vision and action has proven critical for modeling real-world dynamics and achieving coherent manipulation.
However, most existing approaches (li2025unified; du2023learning; zhu2025unified; zhang2025up; cot) adopt a monolithic and goal-agnostic generation strategy. As illustrated in Fig. 1(a), they directly predict the entire rollout of future observations and actions from the current observation and instruction, without explicitly modeling how actions influence observations, or how the target goal should guide the trajectory. This design leads to two fundamental limitations: (1) Goal-Agnostic Observation Generation. Without an explicit goal, models lack semantic guidance during rollout prediction. This often results in visually plausible but task-irrelevant sequences, undermining interpretability and degrading downstream planning accuracy (liu2025efficient; hu2023planning; brohan2023can). (2) Implicit Observation-Action Modeling. Observations and actions are often generated in parallel or via loosely coupled pathways, lacking explicit modeling of their causal interplay. This weakens temporal coherence, limits adaptability, and hinders the mutual reinforcement between perception and manipulation essential for decision-making.
@@ 2  Related Work
Recent works have shown that conditioning for goal states can significantly improve planning (luo2024grounding; bae2024tldr; gong2024goal; rens2025proposing; li2024optimus; li2025optimus; zhang2025falcon; lyu2025puma; shao2019multi; shen2024mome; chen2025SimpAgent). LBP (liu2025efficient) learns the dynamics conditioned by latent goals and performs backward planning from the goal, improving the manipulation in the simulation. In the real world, SayCan (brohan2023can) combines language instructions with value grounding to guide robotic behavior, demonstrating strong performance in instruction-following tasks. Hu et al.(hu2023planning) propose Planning Exploratory Goals (PEG) to encourage intrinsic goal-driven exploration, while Li et al.(li2022hierarchical) decompose long-horizon tasks into goal-conditioned subtasks using hierarchical policies. Although these approaches leverage goal conditioning to improve efficiency or task decomposition, they often treat goal information as a constraint or policy input, without explicitly incorporating it into the perception-action generation process. In contrast, our method introduces a goal observation as a semantic anchor and explicitly integrates it into both observation synthesis and action refinement.
@@
Anticipating future visual observations has proven to be a powerful mechanism for enhancing policy learning in robotics (hu2024video; lu2025h; huang2025enerverse; zhao2024vlmpc; xu2025vilp; zhu2025unified; shao2023detecting). UWM (zhu2025unified) integrates action and video diffusion within a unified transformer, enhancing robustness and generalization in imitation learning. UVA (li2025unified) introduces the unified video action model, which jointly optimizes video and action predictions to achieve high accuracy. UP-VLA (zhang2025up) trains a unified vision-language-action model with joint multimodal understanding and future prediction objectives, enhancing both high-level semantic reasoning and low-level spatial understanding. PAD (guo2024prediction) unifies image prediction and robot action within a joint denoising process. However, these approaches often treat observation and action generation as parallel or loosely coupled processes, lacking structured interaction and goal conditioning. In contrast, our hierarchical coarse-to-fine framework couples observation and action bidirectionally, using goal-grounded prediction and intermediate visual states to enable temporally coherent, semantically aligned manipulation.
@@ 3.1  Preliminaries and Problem Formulation
In robotic manipulation, recent studies (cot; du2023learning; liang2024dreamitate; li2025unified) have demonstrated that predicting future visual observations can significantly enhance policy learning. A common paradigm in such frameworks is defined as follows: Given a task instruction $I$ and a sequence of past visual observations $\{O_{t-h+1},\dots,O_{t}\}$ over a history horizon $h$, the objective is to jointly predict a sequence of future action chunks $\{A_{t},\dots,A_{t+h^{\prime}-1}\}$ and the corresponding future observations $\{O_{t+1},\dots,O_{t+h^{\prime}}\}$ over a future horizon $h^{\prime}$. Each action chunk $A_{t}\in\mathbb{R}^{K\times D}$ encodes $K$ consecutive low-level manipulation steps, where each step is represented as a $D$-dimensional action vector.
@@ 3.2  H-GAR: Overview
Following prior works (chang2022maskgit; li2024autoregressive; li2025unified), we encode each image observation into a sequence of $N$ latent tokens using a pretrained VAE (rombach2022high), followed by a linear projection. Let $\mathbf{V}_{\mathcal{H}}$ and $\mathbf{V}_{\mathcal{F}}$ denote the tokenized representations from past and masked future observations, respectively. The task instruction $I$ is encoded into language embedding $T_{I}$ using a pretrained CLIP text encoder (radford2021learning). All tokens are concatenated channel-wise to form a multimodal sequence, which is then processed by a Transformer encoder to produce joint latent representations for future steps:
@@
To generate a semantic goal anchor, we adopt a lightweight video diffusion decoder (li2024autoregressive) conditioned on the final latent $\mathbf{Z}_{t+h^{\prime}}$. Each token $\mathbf{z}_{i}\in\mathbf{Z}_{t+h^{\prime}}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{N}\}$ predicts a visual patch through a denoising process, which is decoded by a pretrained VAE to reconstruct the goal frame $O_{t+h^{\prime}}$. The training objective minimizes the denoising error:
@@
Diffusion Policy-C* [RSS’23] (diff) 0.91 3 0.68 4 0.53 9
Diffusion Policy-T* [RSS’23] (diff) 0.78 5 0.63 5 0.58 6
UniPi* [NeurIPS’23] (du2023learning) 0.42 6 0.19 7 - -
OpenVLA* [CoRL’24] (openvla) 0.35 7 0.22 6 0.54 8
STAR* [ICML’25] (star) - - - - 0.89 3
CoT-VLA* [CVPR’25] (cotvla) - - - - 0.69 5
SpatialVLA* [RSS’25] (spatialvla) - - - - 0.56 7
PD-VLA† [arXiv’25] (pdvla) 0.82 4 0.71 3 0.92 2
UVA† [RSS’25] (li2025unified) 0.96 2 0.85 2 0.89 3
@@
VQ-BeT* (be) 5/10 3/10 4/10 3/10 1/10 - -
QueST* (quest) 6/10 4/10 3/10 1/10 0/10 - -
STAR* (star) 8/10 6/10 6/10 4/10 3/10 - -
PD-VLA† (pdvla) 8/10 7/10 6/10 6/10 4/10 6/10 4/10
UVA† (li2025unified) 7/10 6/10 6/10 5/10 3/10 5/10 3/10
@@
Table 3: Observation Generation Results. Evaluation on simulated (Libero-10) and real-world (Mouse Arrangement) environments across different autoregressive steps. Lower FVD (unterthiner2018towards) reflects better visual fidelity.
@@ 4.1  Experiment Settings
Our simulation experiments include two settings: Single-Task Evaluation and Multi-Task Evaluation. In the single-task setting, we train and evaluate separate policies for each task, using PushT (chi2023diffusion; florence2022implicit) as a representative example. In the multi-task setting, a single policy is trained to handle multiple task goals. Following standard protocols, we evaluate on PushT-M (li2025unified) and Libero-10 (liu2023libero). Each task is tested in 50 diverse environments with different random seeds. Detailed descriptions of these simulation benchmarks are provided in Appendix B.1.
@@ 5  Conclusion
We provide comprehensive supplementary material to support the methodology, implementation, and analysis introduced in the main paper on H-GAR.
@@ A Implementation Details
Early autoregressive visual modeling approaches, such as VQGAN-based methods (esser2021VQGAN), convert images into discrete token sequences via vector quantization. During training, a subset of tokens is randomly masked, and the model learns to reconstruct the missing content. This setup enables autoregressive generation during inference, where images or videos are synthesized step-by-step by progressively unmasking tokens. While effective, this method often suffers from quantization artifacts and loss of fine-grained details, resulting in blurry or distorted outputs.
To mitigate this limitation, more recent work has shifted towards continuous latent modeling. Instead of predicting discrete indices, these methods directly model continuous latent features using a diffusion process, which allows for richer and smoother reconstructions. For example, Recent work (li2024autoregressive) replaces discrete token modeling with a continuous diffusion-based representation, achieving improved visual quality.
@@ B.3 Baselines
  • Diffusion Policy-C/T (Chi2023diffusion_policy) are two variants of Diffusion Policy, where the “C” version adopts a CNN-based policy backbone, while the “T” version replaces it with a Transformer-based architecture. Both variants generate fine-grained actions based on past observations and are widely adopted in robotic imitation learning benchmarks.
@@
  • OpenVLA (Kim2024openvla) is a vision-language-action model built on a pretrained LLaMA-2 backbone. It supports both image and language modalities and is designed to enable multi-task generalization.
@@
  • SpatialVLA (spatialvla) enhances vision-language-action modeling by incorporating egocentric 3D spatial reasoning. It introduces Ego3D Position Encoding, which integrates depth cues into visual features without needing camera calibration, and Adaptive Action Grids, which discretize actions into spatial tokens.
@@
  • CoT-VLA (cotvla) introduces visual chain-of-thought reasoning by first generating a subgoal image based on the instruction and current observation, then predicting actions to reach it. This two-stage process improves planning and generalization in complex tasks. CoT-VLA is built on VILA-U with a hybrid attention mechanism. It leverages large-scale pretraining on both action-labeled and unlabeled videos.
@@
  • UniPi (unipi) formulates policy learning as text-conditioned video generation. It first synthesizes future observations based on the instruction, then infers actions via an inverse dynamics model. This decoupled formulation promotes generalization and adaptability across tasks.
@@
  • UVA (li2025unified) is a generative model for embodied agents that produces a unified latent representation, which is decoded via separate diffusion heads for visual trajectory and action sequence generation. This design supports joint modeling while maintaining modularity between visual and control pathways.
@@
  • VQ-BeT (vqbet) employs a single-layer MLP encoder with 128 dimensions and a residual vector quantization module containing roughly 1024 codes. The model processes sequences using an observation window of 10 timesteps and an action window size of $T=5$, as longer sequences (e.g., $T=32$) may lead to information degradation due to over-compression.
@@
  • QueST (quest) employs a 6-layer transformer decoder with 6 attention heads and an embedding dimension of 384. It generates actions using a discrete vocabulary of size 1000, with token embeddings matching the model dimension. During inference, beam search is performed with a width of 5 and temperature set to 1.0, predicting over action blocks of 8 timesteps.
@@
  • STAR (star) is a skill-based control framework that incorporates RaRSQ to preserve geometric structure during skill quantization and a Causal Skill Transformer (CST) to autoregressively model skill dependencies and offsets. It shows strong performance on long-horizon manipulation tasks.
@@
  • PD-VLA (pdvla) is a vision-language-action model designed to improve decoding efficiency when integrated with action chunking. Traditional autoregressive VLA models suffer from inference latency due to sequential action token generation. PD-VLA addresses this by reformulating the decoding process into a parallel fixed-point iteration system, enabling all action tokens to be predicted simultaneously in each iteration. This approach accelerates inference without requiring retraining or architectural modifications.

### v2

@@ 1  Introduction
Effective planning and manipulation in robotics (Zhou et al. 2025b; Li et al. 2024a; Zhong et al. 2025; Zhou et al. 2025a; Wu et al. 2023; Bharadhwaj et al. 2024; Zhu et al. 2025b, c; Li et al. 2025b; Xie et al. 2025; Shao et al. 2024; Li et al. 2025c) require the ability to anticipate both how the environment evolves and how actions unfold over time. To this end, a growing body of research (Zhao et al. 2025a; Liang et al. 2024; Li et al. 2025a) has shown that jointly predicting future visual observations and action sequences brings substantial benefits to robotic manipulation. This owes to the fact that anticipated visual observations provide rich contextual cues, including spatial layouts, object affordances, and interaction dynamics, which help guide action planning. In parallel, predicted actions expose the causal structure of interactions, enabling more accurate forecasting of how scenes evolve over time. This bidirectional coupling between vision and action has proven critical for modeling real-world dynamics and achieving coherent manipulation.
However, most existing approaches (Li et al. 2025a; Du et al. 2023b; Zhu et al. 2025a; Zhang et al. 2025a; Zhao et al. 2025a) adopt a monolithic and goal-agnostic generation strategy. As illustrated in Fig. 1(a), they directly predict the entire rollout of future observations and actions from the current observation and instruction, without explicitly modeling how actions influence observations, or how the target goal should guide the trajectory. This design leads to two fundamental limitations: (1) Goal-Agnostic Observation Generation. Without an explicit goal, models lack semantic guidance during rollout prediction. This often results in visually plausible but task-irrelevant sequences, undermining interpretability and degrading downstream planning accuracy (Liu et al. 2025; Hu et al. 2023; Brohan et al. 2023). (2) Implicit Observation-Action Modeling. Observations and actions are often generated in parallel or via loosely coupled pathways, lacking explicit modeling of their causal interplay. This weakens temporal coherence, limits adaptability, and hinders the mutual reinforcement between perception and manipulation essential for decision-making.
@@ 2  Related Work
Recent works have shown that conditioning for goal states can significantly improve planning (Luo and Du 2024; Bae, Park, and Lee 2024; Gong et al. 2024; Rens 2025; Li et al. 2024c, 2025d; Zhang et al. 2025b; Lyu et al. 2025; Shao et al. 2019; Shen et al. 2024; Chen et al. 2025). LBP (Liu et al. 2025) learns the dynamics conditioned by latent goals and performs backward planning from the goal, improving the manipulation in the simulation. In the real world, SayCan (Brohan et al. 2023) combines language instructions with value grounding to guide robotic behavior, demonstrating strong performance in instruction-following tasks. Hu et al.(Hu et al. 2023) propose Planning Exploratory Goals (PEG) to encourage intrinsic goal-driven exploration, while Li et al.(Li et al. 2022) decompose long-horizon tasks into goal-conditioned subtasks using hierarchical policies. Although these approaches leverage goal conditioning to improve efficiency or task decomposition, they often treat goal information as a constraint or policy input, without explicitly incorporating it into the perception-action generation process. In contrast, our method introduces a goal observation as a semantic anchor and explicitly integrates it into both observation synthesis and action refinement.
@@
Anticipating future visual observations has proven to be a powerful mechanism for enhancing policy learning in robotics (Hu et al. 2024; Lu et al. 2025; Huang et al. 2025; Zhao et al. 2024; Xu, Qiu, and She 2025; Zhu et al. 2025a; Shao, Wu, and Liu 2023). UWM (Zhu et al. 2025a) integrates action and video diffusion within a unified transformer, enhancing robustness and generalization in imitation learning. UVA (Li et al. 2025a) introduces the unified video action model, which jointly optimizes video and action predictions to achieve high accuracy. UP-VLA (Zhang et al. 2025a) trains a unified vision-language-action model with joint multimodal understanding and future prediction objectives, enhancing both high-level semantic reasoning and low-level spatial understanding. PAD (Guo et al. 2024) unifies image prediction and robot action within a joint denoising process. However, these approaches often treat observation and action generation as parallel or loosely coupled processes, lacking structured interaction and goal conditioning. In contrast, our hierarchical coarse-to-fine framework couples observation and action bidirectionally, using goal-grounded prediction and intermediate visual states to enable temporally coherent, semantically aligned manipulation.
@@ 3.1  Preliminaries and Problem Formulation
In robotic manipulation, recent studies (Zhao et al. 2025a; Du et al. 2023b; Liang et al. 2024; Li et al. 2025a) have demonstrated that predicting future visual observations can significantly enhance policy learning. A common paradigm in such frameworks is defined as follows: Given a task instruction $I$ and a sequence of past visual observations $\{O_{t-h+1},\dots,O_{t}\}$ over a history horizon $h$, the objective is to jointly predict a sequence of future action chunks $\{A_{t},\dots,A_{t+h^{\prime}-1}\}$ and the corresponding future observations $\{O_{t+1},\dots,O_{t+h^{\prime}}\}$ over a future horizon $h^{\prime}$. Each action chunk $A_{t}\in\mathbb{R}^{K\times D}$ encodes $K$ consecutive low-level manipulation steps, where each step is represented as a $D$-dimensional action vector.
@@ 3.2  H-GAR: Overview
Following prior works (Chang et al. 2022; Li et al. 2024b, 2025a), we encode each image observation into a sequence of $N$ latent tokens using a pretrained VAE (Rombach et al. 2022), followed by a linear projection. Let $\mathbf{V}_{\mathcal{H}}$ and $\mathbf{V}_{\mathcal{F}}$ denote the tokenized representations from past and masked future observations, respectively. The task instruction $I$ is encoded into language embedding $T_{I}$ using a pretrained CLIP text encoder (Radford et al. 2021). All tokens are concatenated channel-wise to form a multimodal sequence, which is then processed by a Transformer encoder to produce joint latent representations for future steps:
@@
To generate a semantic goal anchor, we adopt a lightweight video diffusion decoder (Li et al. 2024b) conditioned on the final latent $\mathbf{Z}_{t+h^{\prime}}$. Each token $\mathbf{z}_{i}\in\mathbf{Z}_{t+h^{\prime}}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{N}\}$ predicts a visual patch through a denoising process, which is decoded by a pretrained VAE to reconstruct the goal frame $O_{t+h^{\prime}}$. The training objective minimizes the denoising error:
@@
Diffusion Policy-C* [RSS’23] (Chi et al. 2023b) 0.91 3 0.68 4 0.53 9
Diffusion Policy-T* [RSS’23] (Chi et al. 2023b) 0.78 5 0.63 5 0.58 6
UniPi* [NeurIPS’23] (Du et al. 2023b) 0.42 6 0.19 7 - -
OpenVLA* [CoRL’24] (Kim et al. 2024b) 0.35 7 0.22 6 0.54 8
STAR* [ICML’25] (Hao et al. 2025) - - - - 0.89 3
CoT-VLA* [CVPR’25] (Zhao et al. 2025b) - - - - 0.69 5
SpatialVLA* [RSS’25] (Qu et al. 2025) - - - - 0.56 7
PD-VLA† [arXiv’25] (Song et al. 2025) 0.82 4 0.71 3 0.92 2
UVA† [RSS’25] (Li et al. 2025a) 0.96 2 0.85 2 0.89 3
@@
VQ-BeT* (Lee et al. 2024a) 5/10 3/10 4/10 3/10 1/10 - -
QueST* (Mete et al. 2024) 6/10 4/10 3/10 1/10 0/10 - -
STAR* (Hao et al. 2025) 8/10 6/10 6/10 4/10 3/10 - -
PD-VLA† (Song et al. 2025) 8/10 7/10 6/10 6/10 4/10 6/10 4/10
UVA† (Li et al. 2025a) 7/10 6/10 6/10 5/10 3/10 5/10 3/10
@@
Table 3: Observation Generation Results. Evaluation on simulated (Libero-10) and real-world (Mouse Arrangement) environments across different autoregressive steps. Lower FVD (Unterthiner et al. 2018) reflects better visual fidelity.
@@ 4.1  Experiment Settings
Our simulation experiments include two settings: Single-Task Evaluation and Multi-Task Evaluation. In the single-task setting, we train and evaluate separate policies for each task, using PushT (Chi et al. 2023c; Florence et al. 2022) as a representative example. In the multi-task setting, a single policy is trained to handle multiple task goals. Following standard protocols, we evaluate on PushT-M (Li et al. 2025a) and Libero-10 (Liu et al. 2023). Each task is tested in 50 diverse environments with different random seeds. Detailed descriptions of these simulation benchmarks are provided in Appendix B.1.
@@ A Implementation Details
Early autoregressive visual modeling approaches, such as VQGAN-based methods (Esser, Rombach, and Ommer 2021), convert images into discrete token sequences via vector quantization. During training, a subset of tokens is randomly masked, and the model learns to reconstruct the missing content. This setup enables autoregressive generation during inference, where images or videos are synthesized step-by-step by progressively unmasking tokens. While effective, this method often suffers from quantization artifacts and loss of fine-grained details, resulting in blurry or distorted outputs.
To mitigate this limitation, more recent work has shifted towards continuous latent modeling. Instead of predicting discrete indices, these methods directly model continuous latent features using a diffusion process, which allows for richer and smoother reconstructions. For example, Recent work (Li et al. 2024b) replaces discrete token modeling with a continuous diffusion-based representation, achieving improved visual quality.
@@ B.3 Baselines
  • Diffusion Policy-C/T (Chi et al. 2023a) are two variants of Diffusion Policy, where the “C” version adopts a CNN-based policy backbone, while the “T” version replaces it with a Transformer-based architecture. Both variants generate fine-grained actions based on past observations and are widely adopted in robotic imitation learning benchmarks.
@@
  • OpenVLA (Kim et al. 2024a) is a vision-language-action model built on a pretrained LLaMA-2 backbone. It supports both image and language modalities and is designed to enable multi-task generalization.
@@
  • SpatialVLA (Qu et al. 2025) enhances vision-language-action modeling by incorporating egocentric 3D spatial reasoning. It introduces Ego3D Position Encoding, which integrates depth cues into visual features without needing camera calibration, and Adaptive Action Grids, which discretize actions into spatial tokens.
@@
  • CoT-VLA (Zhao et al. 2025b) introduces visual chain-of-thought reasoning by first generating a subgoal image based on the instruction and current observation, then predicting actions to reach it. This two-stage process improves planning and generalization in complex tasks. CoT-VLA is built on VILA-U with a hybrid attention mechanism. It leverages large-scale pretraining on both action-labeled and unlabeled videos.
@@
  • UniPi (Du et al. 2023a) formulates policy learning as text-conditioned video generation. It first synthesizes future observations based on the instruction, then infers actions via an inverse dynamics model. This decoupled formulation promotes generalization and adaptability across tasks.
@@
  • UVA (Li et al. 2025a) is a generative model for embodied agents that produces a unified latent representation, which is decoded via separate diffusion heads for visual trajectory and action sequence generation. This design supports joint modeling while maintaining modularity between visual and control pathways.
@@
  • VQ-BeT (Lee et al. 2024b) employs a single-layer MLP encoder with 128 dimensions and a residual vector quantization module containing roughly 1024 codes. The model processes sequences using an observation window of 10 timesteps and an action window size of $T=5$, as longer sequences (e.g., $T=32$) may lead to information degradation due to over-compression.
@@
  • QueST (Mete et al. 2024) employs a 6-layer transformer decoder with 6 attention heads and an embedding dimension of 384. It generates actions using a discrete vocabulary of size 1000, with token embeddings matching the model dimension. During inference, beam search is performed with a width of 5 and temperature set to 1.0, predicting over action blocks of 8 timesteps.
@@
  • STAR (Hao et al. 2025) is a skill-based control framework that incorporates RaRSQ to preserve geometric structure during skill quantization and a Causal Skill Transformer (CST) to autoregressively model skill dependencies and offsets. It shows strong performance on long-horizon manipulation tasks.
@@
  • PD-VLA (Song et al. 2025) is a vision-language-action model designed to improve decoding efficiency when integrated with action chunking. Traditional autoregressive VLA models suffer from inference latency due to sequential action token generation. PD-VLA addresses this by reformulating the decoding process into a parallel fixed-point iteration system, enabling all action tokens to be predicted simultaneously in each iteration. This approach accelerates inference without requiring retraining or architectural modifications.